# Periodic robust robotic rock chop via virtual model control

### v1

@@ Abstract
Robotic cutting is a challenging contact-rich manipulation task where the robot must simultaneously negotiate unknown object mechanics, large contact forces, and precise motion requirements. We introduce a new virtual-model control scheme that enables knife rocking motion for robot manipulators, without pre-planned trajectories or precise information of the environment. Motion is generated through interconnection with virtual mechanisms, given by virtual springs, dampers, and masses arranged in a suitable way. Through analysis and experiments, we demonstrate that the controlled robot behavior settles into a periodic motion. Experiments with a Franka manipulator demonstrate robust cuts with five different vegetables, and sub-millimeter slice accuracy from 1 mm to 6 mm at nearly one cut per second. The same controller survives changes in knife shape and cutting board height, and adaptation to a different humanoid manipulator, demonstrating robustness and platform independence.
Index Terms:
@@
Dexterous Manipulation; Compliance and Impedance Control; Robust/Adaptive Control of Robotic Systems; Motion Control of Manipulators.
I Introduction
As a complex manipulation task, robotic cutting requires effective control of the interaction between the cutting tool and its environment, including the geometries of specific objects and their material properties. Contact-rich manipulation is a central theme of robotics research, with significant progress achieved through learning-based approaches [1, 2, 3, 4, 5] and model-based approaches [6, 7, 8, 9]. At the low level of control algorithms, impedance control [10, 11, 12] is often deployed to tame the complex interaction with the environment. However, despite the significant advances in robotic manipulation, robotic cutting remains an open problem. The complex interaction among the knife, the material being cut, and the surrounding environment makes trajectory planning a non-trivial task. Furthermore, determining how to dynamically adjust the manipulator’s impedance in a task-relevant manner is an ongoing area of research.
@@ I Introduction
Consider the ‘trivial’ task of cutting in a kitchen. The knife repeatedly comes into contact with objects of varying material properties, ranging from soft fruits, such as strawberries, to stiffer objects, such as pumpkins and a rigid cutting board. The soft mechanics of food make accurate simulation challenging and this, in turn, complicates the application of learning- and optimization-based methods. In practical settings, learning approaches require a large number of interactions to converge to an effective control policy [13, 14, 15, 16], which not only produces significant food waste but also poses safety risks when experimenting with a knife. Moreover, successful cutting is not just about separating the food; it must also satisfy additional criteria, such as achieving an optimal cut frequency for efficiency and ensuring precise slice thickness to meet culinary presentation standards [17].
@@
Figure 1: Snapshots from robotic cutting.
@@
Accurate trajectory tracking and planning optimization do not seem to yield satisfactory results in the case of cutting. Consider the rocking chop - a technique commonly used in vegetable cutting where the knife is rolled back and forth. Executing such a motion based on controlled predefined trajectories requires precise knowledge of the blade’s geometry to plan a corresponding trajectory [18]. It also requires an exact knowledge of the environment, namely the height and geometry of the cutting board and perhaps of the food. In contrast, it would be highly desirable to have a robust approach that does not depend on precise information, such as the knife geometry. Conventional trajectory generation methods lack this generalization capability and typically rely on accurate vision sensing for re-planning [19, 20, 21].
In this paper, we propose a novel approach that is less reliant on accurate information. Our approach is a virtual model controller (VMC) [22, 23] capable of generating complex motions for contact-rich manipulation tasks. In our approach, the desired robot behavior emerges from the interaction between the robot’s physical mechanics, a virtual mechanism given by an organized collection of mass, springs, and dampers tailored to the task, and the environment. For cutting, these virtual components can realize the necessary motion without requiring explicit trajectory or force tracking. The cutting action remains feasible even if the knife, cutting board or object changes, though the cutting performance may degrade and the resulting motion varies. Moreover, by incorporating sensor feedback, cutting state can be monitored and virtual components can be optimized online to enhance cutting performance such as frequency and force needed for successful separation. Importantly, our controller does not require extensive training data, and tuning is efficient thanks to the intuitive virtual mechanism design.
The central hypothesis of this work is that virtual model control offers a robust solution for interactive tasks, addressing the inherent limitations of trajectory-based methods, specifically their reliance on precise environmental information and the computational burden of adapting to changes. This paper contains the following contributions:
  • a virtual mechanisms controller that realizes rocking motion without explicit trajectory planning;
  • an analytical and experimental demonstration that the controlled robot produces a periodic cutting motion;
  • demonstration of successful cutting of several vegetables, with variable thicknesses and speeds, analyzed through a selection of performance metrics;
  • demonstration of robust cutting, through testing against different environments;
  • real-time adaptation of the virtual mechanism parameters, to improve performance.
@@
The paper is organized as follows. A brief overview of the robotic cutting landscape is summarized in Section II. The methodology of virtual model control and a detailed description of the virtual model controller for cutting are provided in Section III. Section IV provides the details of the implementation and an illustration of the robot motion. An energy-based analysis of the controlled robot behavior can be found in Section V. Experimental benchmarks are discussed in Section VI, with a focus on performance and robustness. Conclusions and further research directions are explored in Section VII, which also provides a detailed comparison with related works.
II Related works
@@
Predicting the mechanical interaction during food cutting is inherently challenging. One of the reasons is that developing an accurate physical model of foodstuffs is complex. The force required to cut through a food item depends on numerous factors, including ripeness, skin presence, orientation, slice thickness, and the knife’s rake angle and sharpness [24, 25, 26, 27]. Simply applying a high cutting force can yield imprecise cuts or even damage delicate produce such as tomatoes. In earlier studies, cutting force was often treated as an external disturbance, with adaptive controllers employing position and velocity feedback to maintain predetermined trajectories [28]. Impedance control was introduced to regulate cutting force directly, augmented by adaptive laws to accommodate time-varying desired positions and environments [29].
@@
Robotic cutting is commonly approached as a motion tracking problem, often informed by visual sensing. A key challenge lies in formulating the cutting trajectory itself. While basic trajectories can be designed, more sophisticated cutting trajectories are harder to derive. Remarkably, [19] obtained a smoothed path from camera data to separate meat muscle. [20] proposed a cut and slice trajectory adjusted through force sensing, to reduce damage to deformable objects. More complex rolling (or rocking) actions have been addressed by leveraging the knife geometry to maintain a constant contact force [18]. [30] explored learning from demonstration, introducing a coupling term via the Udwadia–Kalaba method to suppress lateral blade movements. Reinforcement learning is used in [16] to learn a cutting trajectory and impedance parameters. Because developing an exact physical model of foodstuffs is challenging, learning-based approaches have also garnered attention. A series of works [13, 14, 15] employs data-driven model predictive control (MPC), where a neural network is used to predict future knife positions based on current position and force measurements, thus enabling real-time optimization of the reference force. In [31], variational autoencoders are used to model the cutting dynamics, though their primary goal is to detect material properties and thickness rather than to learn a cutting strategy.
@@
A second challenge involves managing unanticipated interactions with the environment. Inaccurate camera-based detection of the object position, or variations in the table height, can lead to collisions or excessive forces if the robot strictly follows a fixed cutting path. This often results in safety triggers when the commanded position is physically unachievable (e.g., below the cutting board). To deal with the interaction, Yan et al. [32, 18] adopt a hybrid scheme that switches to impedance or direct force control upon detecting contact with the cutting surface. Their method segments the cutting process into distinct press, push, and slice phases, applying different controllers in each stage. In our approach, the controller does not require the identification of different stages of cutting. Rocking is the result of the continuous interaction between robot mechanics, virtual model controller, and environment. Unanticipated interactions may degrade the performance, but only large perturbations will disrupt the whole rocking motion.
A consistent benchmarking framework remains largely absent in the field of robotic cutting. Although prior work often reports end-effector trajectory and force sensor data for individual cutting trials, these metrics are insufficient for real-world applications. In particular, reliability—whether a robot can repeatedly cut and separate multiple slices with accurate thickness—remains understudied. To address this gap, we quantify the performance of our controller against a set of relevant metrics such as cut frequency, cutting force, cutting speed, slice-thickness accuracy and variance.
III Virtual mechanism design for cutting
III-A Virtual model control in a nutshell
A virtual model controller consists of a (virtual) mechanical linkage combined with springs, dampers, and masses that, once coupled with the robot, shapes its behavior. It achieves this by imposing constraints on the motion of the robot and its reaction to the environment. The associated virtual forces are then realized through the robot’s available actuation, typically mapped into joint torques and executed by the motors at each joint [22, 33, 23]. This mapping relies on the robot’s kinematics.
For fully actuated robot manipulator based on revolute joints, using $\boldsymbol{q}$ to denote the joint coordinates of the robot, the force $\boldsymbol{f}_{i}$ generated by a virtual component at a point $\boldsymbol{h}_{i}(\boldsymbol{q})$ of the robot is mapped into corresponding joint torques $\boldsymbol{\tau}$ using the Jacobian matrix $J_{i}(\boldsymbol{q})=\frac{\partial\boldsymbol{h}_{i}(\boldsymbol{q})}{\partial\boldsymbol{q}}$. From the principle of virtual work [34], the torque command to emulate the forces of the virtual model controller is given by
@@ III-A Virtual model control in a nutshell
Virtual model control is intrinsically stable, because the controlled robot is the interconnection of the passive mechanics of the robot and the passive virtual mechanics [23]. Stability is guaranteed by passivity theory [35, 36, 37, 38, 39, 40, 41]. Virtual model control provides an intuitive, task-oriented method for energy shaping and damping injection in robotics [42, 43].
The selection of spring and damper parameters and their geometric configuration also shapes the impedance of the robot [10, 44]. Although impedance control is typically implemented with linear springs and dampers, a nonlinear spring where force saturates can limit excessively rapid movements and enhance safety. In this work, all springs in the virtual mechanism have a saturating nonlinear characteristic between displacement and force given by
@@
Figure 2:  [^20] N saturated nonlinear spring force profile.
III-B Virtual model control for cutting
@@
(b) Left: Maintaining contact. A spring and damper link the knife tip $\boldsymbol{p}_{1}$ to $\boldsymbol{p}_{a}$, keeping the blade in contact with the board while being able to slide along the rocking direction ($y$ axis). Middle and right: Cutting and raising. Spring-damper pair connects the knife handle $\boldsymbol{p}_{2}$ to $\boldsymbol{p}_{b}$ which moves vertically for cutting and raising the knife.
@@
Positions $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are defined relative to the robot’s end-effector, assuming that the knife grip is rigid enough, so that the relationship between end-effector and knife does not change. $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ lie in the world frame and depend on the position of the food. The difference in $y$ coordinate of $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ regulates the slicing distance. The cutting force is regulated through the stiffness parameters $k_{1}$ and $k_{2}$. The position of $\boldsymbol{p}_{a}$ and the selection of the reference $\boldsymbol{r}_{2,1}$ also affect the cutting force.
The switch from cutting to raising is triggered by the knife’s orientation. When $z$ coordinates of $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are close (within tolerance $\delta_{1}$), the knife spine (the line that passes $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$) aligns with the cutting board surface and the cut is done. The switch from raising to cutting is triggered when $\boldsymbol{p}_{2}$ is close to $\boldsymbol{r}_{2,2}$ (within tolerance $\delta_{2}$ measured along the z coordinates), to ensure enough height to cut the food.
@@
where $\boldsymbol{p}_{b}$ is switched between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ depending on whether it is in the cutting or raising phase. The way $\boldsymbol{p}_{b}$ switches is also modulated by the dynamics in Fig. 3(c). We smooth the transition by placing a mass that slides along the line between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$. The mass is connected to $\boldsymbol{p}_{2}$ and $\boldsymbol{r}_{2}$ with springs and dampers.

### v2

@@ Abstract
Robotic cutting is a challenging contact-rich manipulation task where the robot must simultaneously negotiate unknown object mechanics, large contact forces, and precise motion requirements. We introduce a new active virtual-model control scheme that enables knife rocking motion for robot manipulators, without pre-planned trajectories or precise information of the environment. Motion is generated and controlled through switching virtual coupling with virtual mechanisms, given by virtual springs, dampers, and masses arranged in a suitable way. Through analysis and experiments, we demonstrate that the controlled robot behavior settles into a periodic motion. Experiments with a Franka manipulator demonstrate robust cuts with five different vegetables, and sub-millimeter slice accuracy from 1 mm to 6 mm at nearly one cut per second. The same controller survives changes in knife shape and cutting board height, and adaptation to a different humanoid manipulator, demonstrating robustness and platform independence.
keywords:
@@
Dexterous Manipulation; Compliance and Impedance Control; Robust/Adaptive Control; Motion Control.
1 Introduction
Robotic cutting captures the complexity of contact-rich tasks. Cutting requires dexterous, precise motion that adapt to different surfaces, knife geometries, and food mechanics. Contact-rich manipulation is a central theme of robotics research, with significant progress achieved through learning-based approaches Popov et al. (2017); Rajeswaran et al. (2018); Gupta et al. (2021); Radosavovic et al. (2021); Chen et al. (2022) and model-based approaches Posa et al. (2014); Marcucci et al. (2017); Hogan and Rodriguez (2020); Aydinoglu and Posa (2022). At the level of control algorithms, contact-rich interactions require impedance control Hogan (1985a, b, c). Advanced control over the force-displacement relationship at the end-effector is essential to regulate complex interactions with the environment. However, despite significant progress in robotic manipulation, autonomous cutting remains an open challenge due to the unpredictable nature of material fracture and deformation. In addition, the specific geometries of the knife and cutting surface are often only partially known. This uncertainty in the mechanical coupling between the tool and the workspace prevents the use of rigid, pre-programmed trajectories and requires a controller that can handle unmodeled geometric constraints.
@@ 1 Introduction
In the ‘trivial’ task of cutting in a kitchen, the knife repeatedly comes into contact with objects of varying material properties, ranging from soft fruits, such as strawberries, to stiffer objects, such as pumpkins and a rigid cutting board. The soft mechanics of food make accurate simulation challenging and this, in turn, complicates the application of learning- and optimization-based methods. In practical settings, learning approaches require a large number of interactions to converge to an effective control policy Ian Lenz and Saxena (2015); Mitsioni et al. (2019, 2021); Beltran-Hernandez et al. (2024), which not only produces significant food waste but also poses safety risks when experimenting with a knife.
@@
Figure 1: Robotic arm with knife ridigly attached, and snapshots of robotic cutting experiments.
@@
In addition, for dynamic cutting techniques such as the rocking chop, where the blade rolls across the board, classical trajectory tracking and optimization often fail to yield satisfactory results. Executing such motions via predefined trajectories requires precise kinematic models of the blade’s geometry Mu et al. (2023). Furthermore, it demands an exact characterization of the environmental constraints, specifically the height and planarity of the cutting board, as well as the evolving geometry of the food. In contrast, it would be highly desirable to have a robust approach that does not depend on precise information, such as the knife geometry. Conventional trajectory generation methods lack this generalization capability and typically rely on accurate vision sensing for re-planning Long et al. (2013, 2014); Wright et al. (2024).
Impedance control Hogan (1985a) is the standard paradigm for tasks involving stiff and uncertain contacts. However, identifying a reliable impedance for dynamic cutting remains a significant challenge. The system must be sufficiently stiff to penetrate the food, maintain a rectilinear cutting plane, and ensure accurate rolling kinematics. Conversely, it must remain compliant enough to guarantee reliable contact with the board, and moderate the contact forces and prevent jamming in the presence of geometric uncertainties. Whether these competing requirements can be reconciled with a static, uniform impedance remains an open question. Our experimental findings suggest that impedance must instead be modulated according to the specific geometry and phase of the cutting cycle. However, to the best of our knowledge, a formal theory for phase-dependent or adaptive impedance design in the context of robotic cutting with geometric uncertainties has yet to be established.
In this paper, we propose an approach that reduces dependence on accurate modelling information while enabling the generation of complex motions for contact-rich manipulation. Our framework is rooted in the classical methodology of Virtual Model Control (VMC), originally developed to manage the stiff, complex interactions inherent in robotic locomotion Pratt et al. (2001); Larby and Forni (2025). We introduce a novel virtual model controller, comprising a network of virtual springs, dampers, and linkages, to establish robust rolling kinematics for the knife. Rather than relying on pre-programmed trajectories, the desired behavior emerges from the dynamic interaction between the robot’s physical embodiment, the virtual mechanism, and the environment. Methodologically, our approach diverges from existing VMC literature by incorporating a regulated injection of energy, resulting in a stable, periodic rock chop motion. This is achieved via controlled switching between two virtual mechanical topologies that correspond to the rising and falling phases of the knife.
Our design of the virtual mechanisms for cutting is based on intuitive physical choices, which can be described graphically in a simple way. Our virtual model controller does not require extensive training data, and parameter tuning is efficient. The controller modulates the robot impedance in a complex, nonlinear fashion throughout the cutting cycle, driven by real-time interaction features. By leveraging its virtual components, the virtual mechanisms realize the necessary dynamics without requiring explicit trajectory or force tracking, and demonstrates robustness to environmental variations; the cutting action remains viable even if knife, workspace, or object geometry change, albeit with graceful performance degradation. Finally, the integration of sensory feedback enables online monitoring and optimization of the virtual parameters, facilitating the refinement of cutting frequency and force to ensure successful material separation.
The central finding of this paper is that Virtual Model Control (VMC) provides a robust design framework for complex manipulation, significantly reducing the reliance on high-fidelity modeling and the associated computational overhead. The specific contributions of this work are:
  • A novel VMC architecture for autonomous rocking chop motions without explicit trajectory planning;
  • The introduction of regulated energy injection into the VMC framework to sustain periodic motion via controlled mechanical switching;
  • Analytical and experimental demonstrations verifying the stability and periodicity of the cutting motion;
  • Experimental characterization of cutting performance across various types of vegetables, for different thicknesses and cutting speeds;
  • Experimental validation of the controller’s robustness against perturbations of knife geometry and cutting board height;
  • Demonstration of the controller’s portability (retargeting) through implementation on two distinct robotic platforms: the Franka FR3 and the RT Corp Scirus 17;
  • A preliminary study of the online adaptation capabilities enabled by the proposed VMC design.
@@
The paper is organized as follows. A brief overview of the robotic cutting landscape is summarized in Related works. The methodology of virtual model control and a detailed description of the virtual model controller for cutting are provided in Virtual cutting mechanisms design. An energy-based analysis of the controlled robot behavior can be found in Energy-based analysis. Experimental platform and cutting motion provides the details of the implementation and an illustration of the robot motion. Experimental benchmarks are discussed in Experimental result and discussion, with a focus on performance and robustness. Conclusions and further research directions are explored in Conclusion and future works, which also provides a detailed comparison with related works.
2 Related works
@@
Predicting the mechanical interaction during food cutting is inherently challenging. One of the reasons is that developing an accurate physical model of foodstuffs is complex. The force required to cut through a food item depends on numerous factors, including ripeness, skin presence, orientation, slice thickness, and the knife’s rake angle and sharpness Vincent and Lillford (1991); Khan and Vincent (1991); Atkins (2009b, a). Simply applying a high cutting force can yield imprecise cuts or even damage delicate produce such as tomatoes. In earlier studies, cutting force was often treated as an external disturbance, with adaptive controllers employing position and velocity feedback to maintain predetermined trajectories Zeng and Hemami (1997). Impedance control was introduced to regulate cutting force directly, augmented by adaptive laws to accommodate time-varying desired positions and environments Jung and Hsia (1999).
@@
Robotic cutting is commonly approached as a motion tracking problem, often informed by visual sensing. A key challenge lies in formulating the cutting trajectory itself. While basic trajectories can be designed, more sophisticated cutting trajectories are harder to derive. Remarkably, Long et al. (2013) obtained a smoothed path from camera data to separate meat muscle. Long et al. (2014) proposed a cut and slice trajectory adjusted through force sensing, to reduce damage to deformable objects. More complex rolling (or rocking) actions have been addressed by leveraging the knife geometry to maintain a constant contact force Mu et al. (2023). Straižys et al. (2023) explored learning from demonstration, introducing a coupling term via the Udwadia–Kalaba method to suppress lateral blade movements. Reinforcement learning is used in Beltran-Hernandez et al. (2024) to learn a cutting trajectory and impedance parameters. Because developing an exact physical model of foodstuffs is challenging, learning-based approaches have also garnered attention. A series of works Ian Lenz and Saxena (2015); Mitsioni et al. (2019, 2021) employs data-driven model predictive control (MPC), where a neural network is used to predict future knife positions based on current position and force measurements, thus enabling real-time optimization of the reference force. In Rezaei-Shoshtari et al. (2020), variational autoencoders are used to model the cutting dynamics, though their primary goal is to detect material properties and thickness rather than to learn a cutting strategy.
@@
A second challenge involves managing unanticipated interactions with the environment. Inaccurate camera-based detection of the object position, or variations in the table height, can lead to collisions or excessive forces if the robot strictly follows a fixed cutting path. This often results in safety triggers when the commanded position is physically unachievable (e.g., below the cutting board). To deal with the interaction, Yan et al. Mu et al. (2019, 2023) adopt a hybrid scheme that switches to impedance or direct force control upon detecting contact with the cutting surface. Their method segments the cutting process into distinct press, push, and slice phases, applying different controllers in each stage. In our approach, the controller does not require the identification of different stages of cutting. Rocking is the result of the continuous interaction between robot mechanics, virtual model controller, and environment. Unanticipated interactions may degrade the performance, but only large perturbations will disrupt the whole rocking motion.
A consistent benchmarking framework remains largely absent in the field of robotic cutting. Although prior work often reports end-effector trajectory and force sensor data for individual cutting trials, these metrics appear insufficient for real-world applications. Reliability, that is, the ability to repeatedly produce separated slices of accurate thickness and handle variations in the environment (e.g., knife, board height, or food properties), remains understudied. In contrast, in this paper we quantify the performance of our controller against a broad set of relevant metrics, such as cut frequency, cutting force, cutting speed, slice-thickness accuracy and variance, and we demonstrate successful cuts across different environmental conditions.
3 Virtual cutting mechanisms design
3.1 Virtual model control in a nutshell
A virtual model controller consists of a (virtual) mechanical linkage combined with springs, dampers, and masses that, once coupled with the robot, shapes its behavior. It achieves this by imposing constraints on the motion of the robot and its reaction to the environment. The associated virtual forces are then realized through the robot’s available actuation, typically mapped into joint torques and executed by the motors at each joint Pratt et al. (2001); Zhang et al. (2024); Larby and Forni (2025). This mapping relies on the robot’s kinematics.
For fully actuated robot manipulator based on revolute joints, using $\boldsymbol{q}$ to denote the joint coordinates of the robot, the force $\boldsymbol{f}_{i}$ generated by a virtual component at a point $\boldsymbol{p}_{i}=\boldsymbol{h}_{i}(\boldsymbol{q})$ of the robot is mapped into corresponding joint torques $\boldsymbol{\tau}$ using the Jacobian matrix $J_{i}(\boldsymbol{q})=\frac{\partial\boldsymbol{h}_{i}(\boldsymbol{q})}{\partial\boldsymbol{q}}$. From the principle of virtual work Spong et al. (2005), the torque command to emulate the forces of the virtual model controller is given by
@@ 3.1 Virtual model control in a nutshell
Virtual model control is intrinsically stable, because the controlled robot is the interconnection of the passive mechanics of the robot and the passive virtual mechanics Larby and Forni (2025). Stability is guaranteed by passivity theory Willems (1972); van der Schaft (1999); Secchi et al. (2007); Stramigioli (2015); Folkertsma and Stramigioli (2017); Rashad et al. (2020); Ortega et al. (2021). Virtual model control provides an intuitive, task-oriented method for energy shaping and damping injection in robotics Ortega et al. (2001, 1998).
The selection of spring and damper parameters and their geometric configuration also shape the impedance of the robot Hogan (1985a); Colgate and Hogan (1988). Although impedance control is typically implemented with linear springs and dampers, nonlinear springs featuring force saturation can limit excessively rapid movements and enhance safety. In this work, all springs in the virtual mechanism have a saturating nonlinear characteristic between displacement and force given by
@@
Figure 2: 2020 N saturated nonlinear spring force profile.
3.2 Virtual model control for cutting
@@
(b) Left: Maintaining contact. A spring and damper link the knife tip $\boldsymbol{p}_{1}$ to $\boldsymbol{p}_{a}$, keeping the blade in contact with the board while being able to slide along the rocking direction ($y$ axis). Middle and right: Cutting and raising. Spring-damper pair connects the knife handle $\boldsymbol{p}_{2}$ to $\boldsymbol{p}_{b}$ which moves vertically for cutting and raising the knife. When switch from cutting to raising, the reference points are shifted along $x$ axis by one slice thickness to produce multiple slices.
@@
Positions $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are defined relative to the robot’s end-effector, assuming that the knife grip is rigid enough, so that the relationship between end-effector and knife does not change. $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ do not have to be on the knife. As long as $\boldsymbol{p}_{1}$ is towards the direction of knife tip and $\boldsymbol{p}_{2}$ is in the direction of handle, the motion is feasible. $\boldsymbol{p}_{a}$, $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ lie in the world frame and depend on the position of the food. The difference in $y$ coordinate of $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ regulates the slicing distance. The cutting force is regulated through the stiffness parameters $k_{1}$ and $k_{2}$. The position of $\boldsymbol{p}_{a}$ and the selection of the reference $\boldsymbol{r}_{2,1}$ also affect the cutting force.
The switch from cutting to raising is triggered by the knife’s orientation. When $z$ coordinates of $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$ are close (within tolerance $\delta_{1}$), the knife spine (the line that passes $\boldsymbol{p}_{1}$ and $\boldsymbol{p}_{2}$) aligns with the cutting board surface and the cut is done. The switch from raising to cutting is triggered when $\boldsymbol{p}_{2}$ is close to $\boldsymbol{r}_{2,2}$ (within tolerance $\delta_{2}$ measured along the z coordinates), to ensure enough height to cut the food. When knife is moving fast, tolerances $\delta_{1}$ and $\delta_{2}$ need to be set large enough to prevent excessive raising or rolling of the knife.
@@
where $\boldsymbol{p}_{b}$ is switched between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ depending on whether it is in the cutting or raising phase. The way $\boldsymbol{p}_{b}$ switches is also modulated by the dynamics in Fig. 3(c). We smooth the transition by placing a mass that slides along the line between $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$. The mass is rigidly attached to $\boldsymbol{p}_{b}$ and connected to $\boldsymbol{r}_{2}$ by spring and dampers.